In the game of​ roulette, a wheel consists of 38 slots numbered​ 0, 00,​ 1, 2,..., 36. To play the​ game, a metal ball is spun around the wheel and is allowed to fall into one of the numbered slots. If the number of the slot the ball falls into matches the number you​ selected, you win​ $35; otherwise you lose​ $1. Complete parts ​(a) through ​(g) below.

Construct a probability distribution for the random variable​ X, the winnings of each spin.

x

​P(x)

35

nothing

minus−1

nothing

​(Type integers or decimals rounded to four decimal places as​ needed.)
Answer #1

SOLUTION

Here, the probability of winning is 1/38 = 0.026315789.


Here, the probability of losing is 1-1/38 = 37/38 or 0.973684211
.

Thus,

x P(x)
35 0.026316
-1 0.973684